2x+3y-1=0
whats the slope and y intercept?

Answer by Alan3354(69443)

Do it like this one:
======================================
whats the slope of
2x-6y=18
--------------
Put the eqn in slope-intercept form y = mx + b
That means solve for y.
2x-6y=18
6y = 2x - 18
y = (1/3)x - 3
---
The slope is the coefficient of x.
m = 1/3
===============
The y-intercept is -3 or (0,-3)